# AT 2025abao: the fourth luminous red nova in M 31

### v2

@@
We present photometric and spectroscopic observations of the luminous red nova (LRN) AT 2025abao, the fourth discovered in M 31. The LRN, associated with the AGB star WNTR23bzdiq, was discovered during the fast rise following the minimum phase. It reached the peak at mag ( mag), and then it settled onto a long-duration plateau in the red bands, lasting 70 days, while it was slowly linearly declining in the blue bands. The object showed similarities at peak with the canonical LRNe V838 Monocerotis, V1309 Scorpii, and with the faint and fast-evolving AT 2019zhd, the third LRN in M31, though the later evolution is different. Spectroscopically, AT 2025abao evolved as a canonical LRN: the early spectra present a blue continuum with narrow Balmer lines in emission; at peak, the spectral continuum has cooled to a yellow colour, with a photospheric temperature of 6000 K. Balmer lines have weakened while absorption lines from metals (Fe i, Fe ii, Sc ii, Ba ii, Ti ii) have developed, and in particular broad (FWHM700 km s-1) from the UV Ca ii H&K lines. Medium- and high-resolution spectra reveal a counter-P Cygni absorption profile in H. Finally, late time spectra show an orange continuum (4000-5000 K), a return in strength of the Balmer lines and the formation of molecular absorption bands. AT 2025abao is the rare case of a LRN with detailed archival information regarding the progenitor system. For the first time, we obtained the spectral energy distribution in the infrared of the precursor of a LRN, which is consistent with that of an M giant/AGB. We propose that the dichotomy of light curve behaviour in LRNe (two peaks vs. plateau) can be explained by the extent and H-richness of the common envelope.
@@
supernovae: individual: AT 2025abao, galaxies: individual: M 311 Introduction
@@
Luminous Red Novae (LRNe; Kulkarni et al. 2007) are astrophysical transients likely resulting from the coalescence of two non-degenerate stars due to the loss of systemic angular momentum following the ejection of a common envelope (e.g., Tylenda et al., 2011; Ivanova et al., 2013b, a; Kochanek et al., 2014; Pejcha, 2014; MacLeod et al., 2022). They are a subclass of “interacting gap transients” (IGTs; Pastorello and Fraser, 2019; Cai et al., 2022b; Reguitti et al., 2026). IGTs have absolute magnitudes at maximum in the range mag, and fill the observational gap that separates the brightest novae from the faintest supernovae. In addition, their photometric and spectroscopic evolution is dominated by the effects of the interaction with a circumstellar medium (CSM, see Pastorello et al. 2019 for a review on the phenomenon).
@@
LRNe are identified through a few recurrent observational characteristics. Firstly, they present a light curve with a double peak. The first peak has a short duration (1 week); it is typically very luminous and characterised by a blue colour. This first maximum is followed a few weeks (or even months) later by a long-lasting second peak or a plateau-like phase, which is redder and usually fainter than the first111However, see the case of AT 2011kp (a.k.a. NGC4490-OT2011; Pastorello et al., 2019; Smith et al., 2016), which had a second peak that was more luminous than the first one..
@@
The spectra of LRNe show a remarkable evolution, with three main phases: at the time of the first peak, the spectra display a blue continuum with narrow Balmer emission lines, similar to those of SNe IIn (Fraser, 2020). At the second peak, which sometimes is more akin to a plateau, the continuum cools to that of a yellow/orange star (5000-6000 K). At this phase, the Balmer lines become very weak, while a forest of absorption lines from neutral and singly-ionized metals (such as O, Ca, Fe, Sc, Ba, V, Ti) dominate the spectrum. Finally, in the third stage - after the second peak - the spectral continuum changes again to resemble that of a red giant star, with a photospheric temperature of only 3000-4000 K. On top of it, the narrow H line in emission becomes prominent again, whereas molecular absorption bands from VO and TiO devour the continuum.
@@
For a number of extragalactic LRNe, the progenitor systems were detected in pre-outburst archival images from deep ground-based facilities or space telescopes (Fraser et al., 2011; Mauerhan et al., 2015; Smith et al., 2016; Blagorodnova et al., 2017; Cai et al., 2019; Pastorello et al., 2021b; Blagorodnova et al., 2021; Cai et al., 2022a). From these findings, it was discovered that the typical progenitors of LRNe are yellow hypergiant stars, with masses up to 50 (Guidolin et al., 2026), and are suggested to sit in the Hertzsprung gap of the HR diagram (Tranin et al., 2025; Addison et al., 2022).
@@
Most LRNe are discovered in distant galaxies, hence they are usually faint targets, making challenging their observational campaigns and, consequently, limiting the amount of information we can infer from their study. For this reason, it is important to find and study IGTs in nearby galaxies, especially within the Local Group. A handful of LRNe were observed in the Milky Way: V4332 Sgr (Martini et al., 1999; Tylenda et al., 2005), OGLE-2002-BLG360 (Tylenda et al., 2013), V838 Mon (Munari et al., 2002; Bond et al., 2003; Tylenda, 2005), and V1309 Sco (Mason et al., 2010; Tylenda et al., 2011). Three others were instead observed in M 31: M31-RV (Mould et al., 1990; Boschi and Munari, 2004), M31-LRN2015, (Kurtenkov et al., 2015; Williams et al., 2015; MacLeod et al., 2017; Lipunov et al., 2017; Blagorodnova et al., 2020), and AT 2019zhd (Pastorello et al., 2021a).
@@
In this work, we present the results of our follow-up campaign of AT 2025abao, the fourth event of its type discovered in the Andromeda Galaxy. The proximity of M 31 (only 750 kpc away) allows us to constrain the progenitor parameters, and to provide a rich dataset of the transient event in terms of cadence, time coverage, resolution, and signal. Another paper was published illustrating the pre-outburst evolution of the source that later generated LRN AT 2025abao (labelled with the survey names WNTR23bzdiq and WTP19aalzlk; Karambelkar et al., 2025b, hereafter K25), and predicting a later evolution towards a LRN. In particular, the discovery of AT 2025abao offered the rare opportunity to perform a detailed analysis of its progenitor system in a stage of relative quiescence, and to study in detail its spectro-photometric properties a short time before the outburst onset222This phase was well monitored in photometry only for the Galactic LRN V1309 Sco (Tylenda et al., 2011).. Through the analysis of infrared data, K25 infer that the progenitor of AT 2025abao was an early asymptotic giant branch (AGB) star with a temperature of about 3,500 K, a luminosity of , a radius of 350 , and a mass of . Over a period of about seven years before the LRN outburst, the star increased its luminosity by a factor of three, with a limited colour/temperature evolution, while the spectrum showed properties compatible with those of an M-type star. The global (though slow) luminosity rise of the source was accompanied by fluctuations in its light curve, compatible with the presence of a binary companion. K25 suggested that the stellar brightening was the consequence of the common envelope evolution of a binary system having an AGB star as a primary member. Our paper will analyse the later evolution on the object, from the late pre-LRN phases to the LRN outburst.
@@
The structure of the paper is the following: in Section 2 we present the discovery circumstances of AT 2025abao, and some observational characteristics. The photometric dataset and the light-curve evolution are presented in Section 3, where the transient is also compared to other LRNe objects with similar properties. The spectral sequence is shown and analysed in Section 4, and finally the results are discussed in Section 5.
@@ 2 Discovery
AT 2025abao was officially discovered on 2025 October 19.8196333Universal Time is used throughout the paper. (Modified Julian Date, or MJD = 60967.8196) by the Mobile Astronomical System of Telescope-Robots program (Kechin et al., 2025), at an unfiltered magnitude of 17.3 (Vega mag), and immediately reported to the Transient Name Server444https://www.wis-tns.org/ (TNS). However, the first advisory of a brightening from a known transient was made by Koichi Itagaki on October 17.4798, at an unfiltered magnitude of 17.8 (Vegamag)555http://www.cbat.eps.harvard.edu/unconf/followups/J00384865+4046079.html. An even earlier pre-discovery measurement was provided by Fabregat et al. (2025) on October 14. The LRN appeared in the south-west quadrant of the very nearby Andromeda galaxy (M 31), at celestial coordinates 00:38:48.62, 40:46:07.6 ( West and South of the nucleus), in the outskirts of its spiral disk. The location of the transient is shown in Figure 1. AT 2025abao was correctly classified on 2025 November 2 as a LRN by Taguchi and Maeda (2025a), who noted similarities with another LRN found in M 31, AT 2019zhd (Pastorello et al., 2021a). For the distance to M 31, we assume the same value adopted by Pastorello et al. (2021a) ( mag, Mpc). Its redshift is (Falco et al., 1999). However, we note that the peaks of the spectral lines of AT 2025abao are blueshifted by . The additional radial velocity is likely caused by the rotation curve of the galaxy disk, with the object being located on the side approaching us at 150 km s-1 (Chemin et al., 2009). The Galactic reddening towards AT 2025abao is modest, although not negligible, mag (Schlafly and Finkbeiner, 2011). Given the peripheral location of AT 2025abao in the disk of M 31, additional contribution to the dust extinction is likely negligible. Hereafter, in agreement with K25, we will only account for the dust attenuation within the Milky Way.
@@ 3.1 Observations and data reduction
While K25 presented early photometric data of AT 2025abao, this work complements their light curve with previously unpublished pre-LRN photometric points from the Zwicky Transient Factory (ZTF; Bellm et al., 2019; Graham et al., 2019). The initial phases following the LRN outburst were exclusively monitored by amateur astronomers and all-sky surveys, such as ZTF and the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018). Our multi-band optical and near-infrared (NIR) follow-up campaign began shortly after the spectroscopic classification of the object, and lasted for approximately four months. The Swift space telescope was also triggered to get ultraviolet (UV) photometry at early phases. The facilities used to collect our photometric data are listed in Table 1. The photometric data were reduced using standard procedures with the dedicated ECsnoopy pipeline666ECsnoopy is a package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at https://sngroup.oapd.inaf.it/ecsnoopy.html.. Instrumental magnitudes were determined through the Point Spread Function (PSF)-fitting technique, and calibrated against the catalogue of the Sloan Digital Sky Survey (SDSS) survey (Ahumada et al., 2020). The Swift/UVOT UV data were reduced with the HEASOFT pipeline v. 6.35.2777NASA High Energy Astrophysics Science Archive Research Center – Heasarc 2014., by aperture photometry within a fixed radius of 5”. We also retrieved calibrated forced photometry measurements from the ALeRCE Broker (Förster et al., 2021) for ZTF, and from the public Forced Photometry server888https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021) for ATLAS. Given its apparent brightness, the object was also observed extensively by amateur astronomers999More observations of AT 2025abao during the outburst are presented by Mikolajczyk et al. (2026).; the magnitudes were measured via aperture photometry on their pre-reduced frames (corrected for bias, flat-field, and dark frames), and calibrated their measurements against the AAVSO Photometric All-Sky Survey101010https://www.aavso.org/apass catalogue (Henden, 2019). The final UV, optical (Sloan, Johnson, and ATLAS), and NIR magnitudes are listed in the CDS, while the light curves are plotted in Fig. 2.
3.2 Light curve, and its interpretation
@@ 3.2.1 Pre-outburst phase
The photometric evolution of the transient during the pre-LRN was extensively studied by K25. The light curve of the transient experienced a slow, long-lasting (about seven years) brightening, with some luminosity fluctuations superposed on that global trend. Additional photometric points presented in this paper show that the source reached a relative maximum about one year before the LRN discovery, followed by a poorly-sampled luminosity decline to a pre-outburst minimum (see, also, Sect. 3.4).
The pre-outburst evolution of AT 2025abao, from the slow luminosity rise to a broad peak, the superposed magnitude fluctuations, and the following luminosity decline, closely resembles that observed in V1309 Sco (Tylenda et al., 2011). In both cases, the peculiar light-curve evolution is likely a consequence of unstable mass transfer between the stellar components of a binary system (Roche Lobe overflow; see, e.g., Pejcha, 2014; Pejcha et al., 2016b). This culminates with the formation of a common envelope embedding the stellar system which becomes progressively optically thick (Ivanova et al., 2013b). A slow pre-LRN luminosity rise was also observed in other LRNe discovered in M 31, such as M31-LRN-2015 (Blagorodnova et al., 2020) and AT 2019zhd (Pastorello et al., 2021a), as well as in several extra-galactic LRNe, although with some limitations due to a lower monitoring cadence and the poorer signal-to-noise datasets111111The slow brightenings heralding LRN eruptions were observed for AT 2014ib (SNhunt248; Kankare et al., 2015; Mauerhan et al., 2015), AT 2015dl (M101-2015OT1; Blagorodnova et al., 2017; Goranskij et al., 2016), AT 2020hat (Pastorello et al., 2021b; Reguitti et al., 2026), AT 2021blu (Pastorello et al., 2023) and, more marginally, in AT 2011kp (Pastorello et al., 2019) and AT 2021biy (Cai et al., 2022a)..

### v3

@@
We present photometric and spectroscopic observations of the luminous red nova (LRN) AT 2025abao, the fourth discovered in M 31. The LRN, associated with the asymptotic giant branch (AGB) star WNTR23bzdiq, was discovered during the fast rise following the minimum phase. It reached its peak at mag ( mag), and then it settled onto a long-duration plateau in the red bands, lasting 70 days, while it was slowly linearly declining in the blue bands. At the peak the object showed similarities with the canonical LRNe V838 Monocerotis, V1309 Scorpii, and the faint and fast-evolving AT 2019zhd, which is the third LRN in M31, though the later evolution is different. Spectroscopically, AT 2025abao evolved as a canonical LRN: the early spectra present a blue continuum with narrow Balmer lines in emission; at the peak, the spectral continuum had cooled to a yellow colour, with a photospheric temperature of 6000 K. Balmer lines had weakened, while absorption lines from metals (Fe i, Fe ii, Sc ii, Ba ii, Ti ii) had developed, and they were particularly broad from the UV Ca ii H&K lines. Medium- and high-resolution spectra reveal narrow (50 km s-1) absorption and broad (450 km s-1) emission profiles in the Balmer lines, from a slower wind and a faster outflow, respectively. Finally, late-time spectra show an orange continuum ( K), a return in strength of the Balmer lines and the formation of molecular absorption bands. AT 2025abao is the rare case of an LRN with detailed archival information regarding the progenitor system. For the first time, we obtained the spectral energy distribution in the infrared of the precursor of an LRN, which is consistent with that of an M giant/AGB. We propose that the dichotomy of light-curve behaviour in LRNe (two peaks vs. plateau) can be explained by the extent and H-richness of the common envelope.
@@
Stars: binaries: close – stars: individual: AT 2025abao – stars: winds, outflows - galaxies: individual: M 311 Introduction
@@
Luminous red novae (LRNe; Kulkarni et al. 2007) are astrophysical transients likely resulting from the coalescence of two non-degenerate stars due to the loss of systemic angular momentum following the ejection of a common envelope (e.g. Tylenda et al., 2011; Ivanova et al., 2013b, a; Kochanek et al., 2014; Pejcha, 2014; MacLeod et al., 2022). They are a subclass of ‘interacting gap transients’ (IGTs; Pastorello and Fraser, 2019; Cai et al., 2022b; Reguitti et al., 2026). IGTs have absolute magnitudes at maximum in the mag range and fill the observational gap that separates the brightest novae from the faintest supernovae. In addition, their photometric and spectroscopic evolution is dominated by the effects of the interaction with a circumstellar medium (CSM; see Pastorello et al. 2019 for a review on the phenomenon).
@@
Luminous red novae are identified through a few recurrent observational characteristics. Firstly, they present a light curve with a double peak. The first peak has a short duration (1 week); it is typically very luminous and characterised by a blue colour. This first maximum is followed, a few weeks (or even months) later, by a long-lasting second peak or a plateau-like phase, which is redder and usually fainter than the first.111However, see the case of AT 2011kp (also known as NGC4490-OT2011; Pastorello et al., 2019; Smith et al., 2016), which had a second peak that was more luminous than the first one.
@@
The spectra of LRNe show a remarkable evolution, with three main phases: at the time of the first peak, the spectra display a blue continuum with narrow Balmer emission lines, similar to those of SNe IIn (Fraser, 2020). At the second peak, which sometimes is more akin to a plateau, the continuum cools to that of a yellow or orange star (5000–6000 K). At this phase, the Balmer lines become very weak, while a forest of absorption lines from neutral and singly ionised metals (such as O, Ca, Fe, Sc, Ba, V, and Ti) dominate the spectrum. Finally, in the third stage – after the second peak – the spectral continuum changes again to resemble that of a red giant star, with a photospheric temperature of only 3000–4000 K. On top of it, the narrow Hline in emission becomes prominent again, whereas molecular absorption bands from VO and TiO devour the continuum.
@@
For a number of extragalactic LRNe, the progenitor systems were detected in pre-outburst archival images from deep ground-based facilities or space telescopes (Fraser et al., 2011; Mauerhan et al., 2015; Smith et al., 2016; Blagorodnova et al., 2017; Cai et al., 2019; Pastorello et al., 2021b; Blagorodnova et al., 2021; Cai et al., 2022a). From these findings, it was discovered that the typical progenitors of LRNe are yellow hypergiant stars, with masses up to 50 (Guidolin et al., 2026), and they are suggested to sit in the Hertzsprung gap of the HR diagram (Tranin et al., 2025; Addison et al., 2022).
@@
Most LRNe are discovered in distant galaxies; hence, they are usually faint targets, making their observational campaigns challenging and, consequently, limiting the amount of information we can infer from their study. For this reason, it is important to find and study IGTs in nearby galaxies, especially within the Local Group. A handful of LRNe were observed in the Milky Way (MW): V4332 Sgr (Martini et al., 1999; Tylenda et al., 2005), OGLE-2002-BLG360 (Tylenda et al., 2013), V838 Mon (Munari et al., 2002; Bond et al., 2003; Tylenda, 2005), and V1309 Sco (Mason et al., 2010; Tylenda et al., 2011). Three others were instead observed in M 31: M31-RV (Mould et al., 1990; Boschi and Munari, 2004), M31-LRN2015, (Kurtenkov et al., 2015; Williams et al., 2015; MacLeod et al., 2017; Lipunov et al., 2017; Blagorodnova et al., 2020), and AT 2019zhd (Pastorello et al., 2021a).
@@
In this work, we present the results of our follow-up campaign of AT 2025abao, the fourth event of its type discovered in the Andromeda galaxy. The proximity of M 31 (only 750 kpc away) allowed us to constrain the progenitor parameters, and to provide a rich dataset of the transient event in terms of cadence, time coverage, resolution, and signal. Another paper was published illustrating the pre-outburst evolution of the source that later generated the LRN AT 2025abao (labelled with the survey names WNTR23bzdiq and WTP19aalzlk; Karambelkar et al., 2025, hereafter K25), and predicting a later evolution towards an LRN. In particular, the discovery of AT 2025abao offered the rare opportunity to perform a detailed analysis of its progenitor system in a stage of relative quiescence and to study its spectro-photometric properties a short time before the outburst onset in detail.222This phase was well monitored in photometry only for the Galactic LRN V1309 Sco (Tylenda et al., 2011). Through the analysis of infrared data, K25 inferred that the progenitor of AT 2025abao was an early asymptotic giant branch (AGB) star with a temperature of about 3,500 K, a luminosity of , a radius of 350 , and a mass of . Over a period of about seven years before the LRN outburst, the star increased its luminosity by a factor of three, with a limited colour and/or temperature evolution, while the spectrum showed properties compatible with those of an M-type star. The global (though slow) luminosity rise of the source was accompanied by fluctuations in its light curve, compatible with the presence of a binary companion. K25 suggested that the stellar brightening was the consequence of the common envelope evolution of a binary system having an AGB star as a primary member. This paper analyses the later evolution of the object, from the late pre-LRN phases to the LRN outburst.
@@
The structure of the paper is as follows: in Section 2 we present the discovery circumstances of AT 2025abao and some observational characteristics. The photometric dataset and the light-curve evolution are presented in Section 3, where the transient is also compared to other LRNe objects with similar properties. The spectral sequence is shown and analysed in Section 4, and finally the results are discussed in Section 5.
@@ 2 Discovery
The LRN AT 2025abao was officially discovered on 2025 October 19333Universal time is used throughout the paper., or modified Julian date (MJD) = 60967.8196, by the Mobile Astronomical System of Telescope-Robots program (Kechin et al., 2025), at an unfiltered magnitude of 17.3 (Vega mag), and immediately reported to the Transient Name Server444https://www.wis-tns.org/ (TNS). However, the first sighting of a brightening from a known transient was made by Koichi Itagaki on October 17.4798, at an unfiltered magnitude of 17.8 (Vegamag).555http://www.cbat.eps.harvard.edu/unconf/followups/J00384865+4046079.html An even earlier pre-discovery measurement was provided by Fabregat et al. (2025) on October 14, at 18.6 mag. The LRN appeared in the south-west quadrant of the very nearby Andromeda galaxy (M 31), at celestial coordinates 00:38:48.62, 40:46:07.6 ( west and south of the nucleus), in the outskirts of its spiral disc. The location of the transient is shown in Fig. 1. AT 2025abao was correctly classified on 2025 November 2 as an LRN by Taguchi and Maeda (2025a), who noted similarities with another LRN found in M 31, AT 2019zhd (Pastorello et al., 2021a). For the distance to M 31, we assumed the same value adopted by Pastorello et al. (2021a) ( mag, Mpc). Its redshift is (Falco et al., 1999). However, we note that the peaks of the spectral lines of AT 2025abao are blueshifted by , which we adopt hereafter. The additional radial velocity is likely caused by the rotation curve of the galaxy disc, with the object being located on the side approaching us at 150 km s-1 (Chemin et al., 2009). The Galactic reddening towards AT 2025abao is modest, although not negligible: mag (Schlafly and Finkbeiner, 2011). Given the peripheral location of AT 2025abao in the disc of M 31, any additional contribution to the dust extinction is likely negligible. Hereafter, in agreement with K25, we only account for the dust attenuation within the MW.
@@ 3.1 Observations and data reduction
While K25 presented early photometric data of AT 2025abao, this work complements their light curve with previously unpublished pre-LRN photometric points from the Zwicky Transient Facility (ZTF; Bellm et al., 2019; Graham et al., 2019). The initial phases following the LRN outburst were exclusively monitored by amateur astronomers and all-sky surveys, such as ZTF and the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018). Our multi-band optical and near-infrared (NIR) follow-up campaign began shortly after the spectroscopic classification of the object and lasted for approximately four months. The Swift space telescope was also triggered to obtain ultraviolet (UV) photometry at early phases. The facilities used to collect our photometric data are listed in Table 2. The photometric data were reduced using standard procedures with the dedicated ECsnoopy pipeline.666ECsnoopy is a package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at https://sngroup.oapd.inaf.it/ecsnoopy.html. Instrumental magnitudes were determined through the point spread function (PSF) fitting technique, and calibrated against the catalogue of the Sloan Digital Sky Survey (SDSS) (Ahumada et al., 2020). The Swift/UVOT UV data were reduced with the HEASOFT pipeline v. 6.35.2777NASA High Energy Astrophysics Science Archive Research Center – Heasarc 2014. by aperture photometry within a fixed radius of 5”. We also retrieved calibrated forced photometry measurements from the ALeRCE Broker (Förster et al., 2021) for ZTF and from the public Forced Photometry server888https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021) for ATLAS. Given its apparent brightness, the object was also observed extensively by amateur astronomers;999More observations of AT 2025abao during the outburst are presented by Mikolajczyk et al. (2026). the magnitudes were measured via aperture photometry on their pre-reduced frames (corrected for bias, flat-field, and dark frames), and calibrated their measurements against the AAVSO Photometric All-Sky Survey101010https://www.aavso.org/apass catalogue (Henden, 2019). The final UV, optical (Sloan, Johnson, and ATLAS), and NIR magnitudes are listed in the CDS, while the light curves are plotted in Fig. 2.
3.2 Light curve and its interpretation
@@ 3.2.1 Pre-outburst phase
The photometric evolution of the transient during the pre-LRN was extensively studied by K25. The light curve of the transient experienced a slow, long-lasting (about seven years) brightening, with some luminosity fluctuations superposed on that global trend. Additional photometric points presented in this paper show that the source reached a relative maximum about one year before the LRN discovery, followed by a poorly-sampled luminosity decline to a pre-outburst minimum (see also Sect. 3.4).
The pre-outburst evolution of AT 2025abao, from the slow luminosity rise to a broad peak, superposed magnitude fluctuations, and ensuing luminosity decline, closely resembles that observed in V1309 Sco (Tylenda et al., 2011). In both cases, the peculiar light-curve evolution is likely a consequence of unstable mass transfer between the stellar components of a binary system (Roche-lobe overflow; e.g. Pejcha, 2014; Pejcha et al., 2016b). This culminates with the formation of a common envelope embedding the stellar system, which becomes progressively optically thick (Ivanova et al., 2013b). A slow pre-LRN luminosity rise was also observed in other LRNe discovered in M 31, such as M31-LRN-2015 (Blagorodnova et al., 2020) and AT 2019zhd (Pastorello et al., 2021a), as well as in several extragalactic LRNe, although with some limitations due to a lower monitoring cadence and the poorer signal-to-noise datasets.111111The slow brightenings heralding LRN eruptions were observed for AT 2014ib (SNhunt248; Kankare et al., 2015; Mauerhan et al., 2015), AT 2015dl (M101-2015OT1; Blagorodnova et al., 2017; Goranskij et al., 2016), AT 2020hat (Pastorello et al., 2021b; Reguitti et al., 2026), AT 2021blu (Pastorello et al., 2023) and, more marginally, in AT 2011kp (Pastorello et al., 2019) and AT 2021biy (Cai et al., 2022a).